# Spectrality of Prime Size Tiles

### v6

@@ Abstract
We prove that if a tile in $\mathbb{Z}^{d}$ has prime size $p$, then it must be spectral. The proof is by contradiction, it is simply shown that the tiling complement of such a tile can not annihilate all $p$-subgroups. In addition, with a simple transformation we prove that any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
@@ 1 Introduction
Let $G$ be a finitely generated Abelian group, a set $A$ is tiling (or is a tile) in $G$ if there exists some $B$ such that $\{A+b\}_{b\in B}$ forms a partition of $G$, and is spectral if there exists a set of group characters of $G$ that forms an orthogonal basis on $L^{2}(A)$. In such cases $B$ (resp. $S$) is said to be a tiling complement (resp. spectrum) of $A$, while $(A,B)$ (resp. $(A,S)$) is called a tiling pair (resp. spectral pair) in $G$, and we often write $G=A\oplus B$.
@@
Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$. The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction. Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit such a spectrum.
@@ 2 Preliminaries
Let $A$ be a multiset (i.e., a set that allows repetitive elements) on $\mathbb{Z}_{n}^{d}$ (i.e., each element of $A$ is a member of $\mathbb{Z}_{n}^{d}$), the multiplicity of an element $a\in A$ (i.e., the number of copies of $a$ in $A$) will be denoted by $\operatorname{mul}(a)$, the characteristic function on such a multiset $A$ will be defined and written as
$\mathbf{1}_{A}(x)=\begin{cases}\operatorname{mul}(x)&x\in A,\\ 0&x\notin A.\end{cases}$
The Fourier transform of $\mathbf{1}_{A}(x)$ on $\mathbb{Z}_{n}^{d}$ is
$\widehat{\mathbf{1}_{A}}(\xi)=\sum_{a\in A}\operatorname{mul}(a)\cdot e^{2\pi i\langle a,\xi\rangle/n},$
where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straightforward to verify that
[…]
where the difference set $\Delta S$ of a usual set (i.e., a set with distinct elements only) $S$ and the zero set $Z(f)$ of a function $f$ is defined respectively as
$\Delta S=\{s-s^{\prime}:s,s^{\prime}\in S,\;s\neq s^{\prime}\},\qquad Z(f)=\{x:f(x)=0\}.$
@@
is its orthogonal group in $\mathbb{Z}_{n}^{d}$. That (3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, then its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
Let the support of a function $f$ in $\mathbb{Z}_{n}^{d}$ be defined as
$\operatorname{supp}(f)=\mathbb{Z}_{n}^{d}\setminus Z(f),$
then the uncertainty principle on finite Abelian groups [3, 20] asserts that
holds for any non-zero function $f$ defined on $\mathbb{Z}_{n}^{d}$.
@@
Let $A$ be a multiset on $\mathbb{Z}_{n}^{d}$, if $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$, then $\mathbf{1}_{A}$ must be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, i.e., $\mathbf{1}_{A}=m\mathbf{1}_{\mathbb{Z}_{n}^{d}}$ for some $m\in\mathbb{N}$.
@@
$Z(\widehat{\mathbf{1}_{\mathbb{Z}_{n}^{d}}})=\mathbb{Z}_{n}^{d}\setminus\{0\}=\Delta\mathbb{Z}_{n}^{d},$
i.e., $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{\mathbb{Z}_{n}^{d}})$, thus combined with the assumption that $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$ we get
@@
so that $\pi_{n}$ is the projection from $\mathbb{Z}^{d}$ to $\mathbb{Z}_{n}^{d}$ (as a map between sets). Given a set $A\subseteq\mathbb{Z}^{d}$, $\pi_{n}(A)$ will denote the multiset obtained by applying $\pi_{n}$ on every element of $A$. Clearly $\pi_{n}(A)$ will only be a usual set if (and only if) the map $A\mapsto\pi_{n}(A)$ is injective.
@@
Let $n=p^{k}m$ with $p$ being a prime number that does not divide $m$, and set $A^{\prime}=\pi_{p^{k}}(A)$. Let both $f,g$ be characteristic functions on $A^{\prime}$ but with the domain of $f$ being the entire $\mathbb{Z}_{n}^{d}$ while the domain of $g$ being $\mathbb{Z}_{p^{k}}^{d}$ only. If $x\in\mathbb{Z}_{n}^{d}$ is an element of order $p^{t}$ for some $1\leq t\leq k$ , then we have
@@
where $x^{\prime}\in\mathbb{Z}_{p^{k}}^{d}$ is the element that satisfies $x=mx^{\prime}$, and the last Fourier transform $\hat{g}$ is performed in $\mathbb{Z}_{p^{k}}^{d}$ only (the first two Fourier transforms $\widehat{\mathbf{1}_{A}}$ and $\widehat{f}$ take place in $\mathbb{Z}_{n}^{d}$).
@@
The support set of $f,g$ coincide, but they are defined on different subgroups, thus we have
$\widehat{f}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/n},\qquad\widehat{g}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/p^{k}}$
in which Fourier transforms in $\mathbb{Z}_{n}^{d}$ and in $\mathbb{Z}_{p^{k}}^{d}$ are applied respectively.
@@
Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
@@
$\widehat{1_{A}}(x)=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle a,x^{\prime}\rangle/m}.$
@@
$\widehat{1_{A}}(y)=\sum_{a\in A}e^{2\pi i\langle a,\lambda x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x^{\prime}\rangle/m}=P(\omega^{\lambda})=0,$
@@ 3 Main results
[…]
@@
[…]
@@
Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
@@
Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ exists since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent). Further set
@@
By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$. This implies that $A$ is also tiling in $\mathbb{Z}^{d}$. Its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$. ∎
Theorem 2 produces some interesting consequences, for example we can assert that any $3$ points must be tiling as long as they are not on a line. On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$).
@@
The author would like to thank Tao Zhang (XDU) for the hospitality in Xi’an, and the Chern Institute of Mathematics (Tianjin) for hosting while preparing for this paper.
The author thanks Jia Zheng (WHU) for mentioning the phenomenon described in Theorem 2 at $p=3$ to him, as well as Shilei Fan (CCNU) and Gergely Kiss (ARIM) for short communications that improved Lemma 5.

### v7

@@ Abstract
We prove that if a tile in $\mathbb{Z}^{d}$ has prime size $p$, then it must be spectral. The proof is by contradiction, it is simply shown that the tiling complement of such a tile can not annihilate all $p$-subgroups. In addition, with a simple transformation we prove that any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
@@ 1 Introduction
Let $G$ be a finitely generated Abelian group, a set $A$ is tiling (or is a tile) in $G$ if there exists some $B$ such that $\{A+b\}_{b\in B}$ forms a partition of $G$, and is spectral if there exists a set $S$ of group characters of $G$ that forms an orthogonal basis on $L^{2}(A)$. In such cases $B$ (resp. $S$) is said to be a tiling complement (resp. spectrum) of $A$, while $(A,B)$ (resp. $(A,S)$) is called a tiling pair (resp. spectral pair) in $G$, and we often write $G=A\oplus B$.
@@
If $p$ is a prime number, then any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$, therefore this necessarily requires $d\geq p-1$. The second theorem leads to some interesting findings, for example, it follows that any $3$ points in $\mathbb{Z}^{d}(d>1)$ must be both spectral and tiling if they are not on the same line. It is also worth mentioning that $p$ linearly independent points must be tiling is already included in [18, Corollary 2.4].
The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$. The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, from each equivalence class we can squeeze out a derived set of size $p$ whose difference set is still in the same class. The existence of a spectrum within one of these classes is then shown by contradiction: Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit a spectrum.
@@ 2 Preliminaries
Let $f$ be a function on $\mathbb{Z}_{n}^{d}$, then its Fourier transform on $\mathbb{Z}_{n}^{d}$ is
$\hat{f}(\xi)=\sum_{x\in\mathbb{Z}_{n}^{d}}f(x)\cdot e^{2\pi i\langle x,\xi\rangle/n},\qquad\xi\in\mathbb{Z}_{n}^{d}.$
Denote by
$\operatorname{supp}(f)=\{x\in\mathbb{Z}_{n}^{d}:f(x)\neq 0\},\qquad Z(f)=\{x\in\mathbb{Z}_{n}^{d}:f(x)=0\},$
the support and the zero set of $f$ respectively, then the uncertainty principle on finite Abelian groups [3, 20] asserts that
holds for any non-zero function $f$ defined on $\mathbb{Z}_{n}^{d}$.
If $A$ is a finite multiset (i.e., a set that allows repetitive elements) on $\mathbb{Z}_{n}^{d}$ (i.e., each element of $A$ is a member of $\mathbb{Z}_{n}^{d}$), and the multiplicity of an element $a\in A$ (i.e., the number of copies of $a$ in $A$) is denoted by $m_{a}$, then the characteristic function on such a multiset $A$ can be defined and written as
$\mathbf{1}_{A}(x)=\begin{cases}m_{x}&x\in A,\\ 0&x\notin A.\end{cases}$
Its Fourier transform on $\mathbb{Z}_{n}^{d}$ is
$\widehat{\mathbf{1}}_{A}(\xi)=\sum_{a\in A}m_{a}\cdot e^{2\pi i\langle a,\xi\rangle/n}.$
If $A,B$ are finite multisets, and $A+B$ is the multiset obtained by enumerating $a\in A,b\in B$ (so $a$ appears $m_{a}$ times and $b$ appears $m_{b}$ times) and take the collection of all sums $a+b$, then straightforward computation shows
$\widehat{\mathbf{1}}_{A+B}=\widehat{\mathbf{1}}_{A}\cdot\widehat{\mathbf{1}}_{B}.$
With these settings we easily get that when $A,B,S$ are usual sets (every element appear only once), then
$(A,S)\text{ is a spectral pair on }\mathbb{Z}_{n}^{d}$ $\qquad\Leftrightarrow\qquad\begin{cases}|A|=|S|,\\ \Delta S\subseteq Z(\widehat{\mathbf{1}}_{A}),\end{cases}$ (2)
$(A,B)\text{ is a tiling pair on }\mathbb{Z}_{n}^{d}$ $\qquad\Leftrightarrow\qquad\begin{cases}|A|\cdot|B|=n^{d},\\ \Delta\mathbb{Z}_{n}^{d}\subseteq Z(\widehat{\mathbf{1}}_{A})\cup Z(\widehat{\mathbf{1}}_{B}),\end{cases}$ (3)
where the difference set $\Delta S$ is defined as
$\Delta S=\{s-s^{\prime}:s,s^{\prime}\in S,\;s\neq s^{\prime}\}.$
@@
is its orthogonal group in $\mathbb{Z}_{n}^{d}$. (4) can be first established on cyclic subgroups (which is quite trivial), then since non-cyclic subgroups are products of cyclic ones, its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
@@
Let $A$ be a multiset on $\mathbb{Z}_{n}^{d}$, if $\mathbb{Z}_{n}^{d}\setminus\{0\}\subseteq Z(\widehat{\mathbf{1}}_{A})$, then $\mathbf{1}_{A}$ must be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, i.e., $\mathbf{1}_{A}=m\mathbf{1}_{\mathbb{Z}_{n}^{d}}$ for some $m\in\mathbb{N}$.
@@
$Z(\widehat{\mathbf{1}}_{\mathbb{Z}_{n}^{d}})=\mathbb{Z}_{n}^{d}\setminus\{0\},$
thus combined with the assumption that $\mathbb{Z}_{n}^{d}\setminus\{0\}\subseteq Z(\widehat{\mathbf{1}}_{A})$ we get
@@
so that $\pi_{n}$ is the projection from $\mathbb{Z}^{d}$ to $\mathbb{Z}_{n}^{d}$ (as a map between sets). Given a finite set $A\subseteq\mathbb{Z}^{d}$, $\pi_{n}(A)$ will denote the multiset obtained by applying $\pi_{n}$ on every element of $A$. Clearly $\pi_{n}(A)$ will only be a usual set if (and only if) the map $A\mapsto\pi_{n}(A)$ is injective.
@@
Let $n=p^{k}m$ with $p$ being a prime number that does not divide $m$, and $A\subseteq\mathbb{Z}_{n}^{d}$. Set $A^{\prime}=\pi_{p^{k}}(A)$, and let $f,g$ both be characteristic functions on $A^{\prime}$, but defined on $\mathbb{Z}_{n}^{d}$ and $\mathbb{Z}_{p^{k}}^{d}$ respectively so that the Fourier transforms
$\widehat{\mathbf{1}}_{A}(x)=\sum_{a\in A}e^{2\pi i\langle a,\xi\rangle/n},\qquad\hat{f}(\xi)=\sum_{a^{\prime}\in A^{\prime}}m_{a^{\prime}}e^{2\pi i\langle a^{\prime},\xi\rangle/n},\qquad\xi\in\mathbb{Z}_{n}^{d},$
take place in $\mathbb{Z}_{n}^{d}$, while
$\hat{g}(\xi^{\prime})=\sum_{a^{\prime}\in A^{\prime}}m_{a^{\prime}}e^{2\pi i\langle a^{\prime},\xi^{\prime}\rangle/p^{k}},\qquad\xi^{\prime}\in\mathbb{Z}_{p^{k}}^{d},$
is performed in $\mathbb{Z}_{p^{k}}^{d}$. If $x=(x_{1},\ldots,x_{d})\in\mathbb{Z}_{n}^{d}$ is an element of order $p^{t}$ for some $1\leq t\leq k$ (which means every coordinate $x_{1},\ldots,x_{d}$ must be divisible by $m$), then we have
@@
where $x^{\prime}=(x_{1}/m,\ldots,x_{d}/m)\in\mathbb{Z}_{p^{k}}^{d}$.
@@
Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $Z(\widehat{\mathbf{1}}_{A})\cap E\neq\emptyset$, then $E$ is completely contained in $Z(\widehat{\mathbf{1}}_{A})$, i.e., $E\subseteq Z(\widehat{\mathbf{1}}_{A})$.
@@
$\widehat{\mathbf{1}}_{A}(x)=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle a,x^{\prime}\rangle/m}.$
@@
$\widehat{\mathbf{1}}_{A}(y)=\sum_{a\in A}e^{2\pi i\langle a,\lambda x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x^{\prime}\rangle/m}=P(\omega^{\lambda})=0,$
@@ 3 Main results
Consider all equivalence classes in $\mathbb{Z}_{n}^{d}$ whose orders are powers of $p$, by Lemma 4, each equivalence class must be either completely inside $Z(\widehat{\mathbf{1}}_{A^{\prime}})$ or completely disjoint with $Z(\widehat{\mathbf{1}}_{A^{\prime}})$.
@@
$\Delta\mathbb{Z}_{p^{k}}^{d}\subseteq Z(\widehat{\mathbf{1}}_{\pi_{p^{k}}(B^{\prime})}).$ (6)
@@
If $p$ is a prime number, then any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
@@
Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ is a group isomorphism between $\mathbb{Z}^{p-1}$ and $V\mathbb{Z}^{p-1}$ since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent). Further set
@@
By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the sublattice $V\mathbb{Z}^{p-1}$. This implies that $A$ is also tiling in $\mathbb{Z}^{d}$. Its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$. ∎
Theorem 2 produces some interesting consequences, for example we can assert that any $3$ points must be tiling as long as they are not on a line. On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (see also [19] for a characterization of tiles with prime power sizes in $\mathbb{Z}$).
@@
The author would like to thank Tao Zhang (XDU) for the hospitality in Xi’an, and the Chern Institute of Mathematics (Tianjin) for hosting while preparing for this paper. The author thanks Jia Zheng (WHU) for mentioning the phenomenon described in Theorem 2 at $p=3$ to him, as well as Shilei Fan (CCNU) and Gergely Kiss (ARIM) for short communications that improved Lemma 5.